OpenStudy (anonymous):

How do I solve (1.5 x 10^-2)/(3 x 10^-4)

OpenStudy (anonymous):

\[\frac{ (1.5 x 10^-2) }{ (3 x 10^-4) }\]

OpenStudy (anonymous):

first you got the the whole thing upside down

OpenStudy (anonymous):

How so? that's how the question is in the textbook.

OpenStudy (anonymous):

how you solve it is different :)

OpenStudy (anonymous):

you cannont have - exponents

OpenStudy (anonymous):

flip it around

OpenStudy (anonymous):

I'm confused.

OpenStudy (anonymous):

(3 x 10^4)/(1.5 x 10^2)

OpenStudy (anonymous):

exponents are now positive

OpenStudy (anonymous):

do you know how to use scientific notation?

OpenStudy (anonymous):

Wait. so if the textbook shows a negative exponent, I just make it positive?

OpenStudy (anonymous):

Oh I see what you did there.

OpenStudy (anonymous):

I got 2 x 10^2 Book shows 5.0 x 10^1

OpenStudy (campbell_st):

well that makes more sense 1.5/3 = 0.5 10^(-2 + 4) = 10^2 so 0.5 x 10^2 = 5 x 10^1

OpenStudy (anonymous):

Well now I'm just even more confused :c...

OpenStudy (anonymous):

3 x 10^4 calculate this = amount ------------ ---------- = answer 1.5 x 10^2 calculate this = amount

OpenStudy (campbell_st):

ok... here is what you need... when dividing the same base subtract the powers \[\frac{x^a}{x^b} = x^{a - b}\] is is what you need for the powers of 10 \[\frac{10^{-2}}{10^{-4}} = 10^{-2 - -4} = 10^2\] then the actual number part is \[\frac{1.5}{3} = 0.5\] so combining this you get the answer \[0.5 \times 10^2\] now in scientific notation you need a number between 1 and 10... so 0.5 isn't in scientific notation. so multiply it by 10 \[10^2 = 10^1 \times 10^1\] so your answer is \[0.5 \times 10^1 \times 10^1 = 5 \times 10^1\]

OpenStudy (anonymous):

Thank you campbell, I get it now.